# Settling dynamics of non-Brownian suspension of spherical and cubic particles in Stokes flow

### v1

@@ 3.4 Probability density function of particle sedimentation velocity
To better understand the differences in velocity fluctuations between spherical and cubic particles, we investigate the instantaneous particle velocity during sedimentation. Figs 6a and 6b illustrate the normalized probability density function (PDF) of particle velocity as a function of bulk concentration. Notably, $U_{i}/U_{\text{sed}}>1$ indicates that particles are falling faster than the average sedimentation velocity. Positive values of $U_{i}/U_{\text{sed}}$ imply downward movement (i.e. in the direction of gravity), while negative values represent upward movement. As the particles start to settle, the surrounding fluid generates an upward jet due to the conservation of mass. In a very dilute regime ($\phi\approx 0.001$), both the spherical and cubic suspensions exhibit a sharp peak, indicating that the velocity variance is small. In this range, the larger inter-particle separation distance results in weaker hydrodynamic interactions, thus deviation from the single particle dynamics are small. Additionally, the percentage of upward-moving particles ($N_{U_{z}}$) is calculated by integrating the area under the PDF of the settling velocity. Figure 6c illustrates that, in the highly dilute regime, the fraction of upward-moving particles approaches to nearly zero for both types of particles.
@@
As the volume fraction increases, velocity fluctuations increase, causing particles to disperse more and collide more frequently. At moderate concentrations, hydrodynamic interactions among nearby settling particles and inter-particle collisions create a strong upward jet, which can lift some particles (both spheres and cubes) along with it. The frequent hydrodynamic interactions and particle rebounds after collisions may also lead to individual particles moving upward. A closer examination of the PDF of vertical particle velocity reveals that the tails become stretched as $\phi$ increases from $0.025$ to $0.125$, with the left tail slightly longer than the right. Therefore, the fraction of upward-moving particles increases with rising bulk concentration in the low to moderate range. Results in Fig. 6c show that in cubic suspensions, $5.1\%$ of particles tend to move against the settling direction, compared to only $4.5\%$ for spherical particles at $\phi\approx 0.1$. The sharp edges and rapid changes in hydrodynamic forces due to variations in orientation make cubes more susceptible to strong translational and rotational displacements, as well as velocity fluctuations. Additionally, cubes tend to acquire slightly higher angular velocities than spheres, resulting in significant rotation-induced forces perpendicular to their direction of movement. Consequently, cubic suspensions generate larger rotation-induced forces due to the higher magnitude of angular rotation, resulting in a greater fraction of particles being entrained by the upward jet compared to spherical suspensions. This phenomenon becomes even more pronounced in the range of $0.250.1<\phi<$ 0.25, where at least $5.8\%$ of cubic particles exhibit upward movement, while only $5.05\%$ of spherical particles do so at $\phi\approx 0.2$ (Fig. 6c). However, further increases in $\phi$ lead to more densely packet system which suppress the velocity fluctuations in both directions. Consequently, the tails of the PDF begin to shrink for both types of particles (Fig. 6a,b), and a gradual decrease in the percentage of upward-moving particles is observed in Fig. 6c.
@@ 3.5 Suspension microstructure
For dilute suspensions ($\phi\simeq 2.5\%$), the RDF reveals a subtle peak for both spherical and cubic particles, though the peak is slightly lower for cubic suspensions (Fig. 7). In this regime, weak hydrodynamic interactions allow particles to form transient pairs, remaining in contact for extended periods due to the faint wakes trailing behind the leading particles. As $\phi$ increases, these wake structures begin to interact with multiple particles, leading to the emergence of aggregates or clusters in close contact, as can be noticed by the appearance of sharper peak in RDF. For cubic particles, the RDF profile appears to be slightly oblique (Fig. 7b), indicating that these particles can interact over a range of center-to-center distances around $r/R=1$, unlike spherical particles, which have a fixed point of contact at $r/R=2$ (Fig. 7a). Thus, at a moderate solid volume fraction ($\phi\simeq 36\%$), we observe a gradual transition in particle-pair distribution: an initial enrichment in particle pairs is evident around $r/R=2.2$ for spheres and $r/R=1.5$ for cubes, which shifts to a depletion of pairs at greater distances from the reference particle, approximately $r/R\simeq 3.5$ for spheres and $r/R\simeq 2.1$ for cubes. Due to the reduced inter-particle spacing compared to dilute conditions, there is an increased tendency for cluster formation at closer distances, a phenomenon more pronounced for spheres, as evidenced by their higher RDF peaks. Cubic particles, however, exhibit a greater tendency to escape clustering due to their higher rotational velocities, a behavior also highlighted in the work of Seyed-Ahmadi and Wachs [37]. With further increases in $\phi$, intensified hindered settling arises from diminished inter-particle distances, resulting in a relatively isotropic particle structure, with persistent oscillations beyond the first RDF peak observed for both suspensions at $\phi\simeq 0.5$. At larger inter-particle distances (e.g., $r/R\geq 5.1$ for spheres and $r/R\geq 3.15$ for cubes), the particle distribution becomes random, with $g(r)\simeq 1$.
@@
As the volume fraction increases, the wakes behind leading particles are increasingly disturbed by nearby particles, leading to more particle pairs clustering around the reference particle and promoting ordering. We observe a transition from a preferential horizontal alignment at lower bulk concentrations ($0.025\leq\phi\leq 0.125$) to a more concentric, ring-like distribution at moderate concentrations ($0.15\leq\phi\leq 0.4$). A spherical contour of particle pairs forms at distances of $r/R=2.05$ to $2.65$ for spheres and $r/R=1.15$ to $1.6$ for cubes at $\phi=0.36$ (Fig. 8) reveling an isotropic ordering. At this volume fraction a second ring due to second neighbors becomes visible.
The most intriguing case occurs at the highest volume fraction studied, $\phi=0.5$, where the pair distribution functions of cubes and spheres exhibit striking differences. For spheres, the pair distribution function shows a high concentration of particles at $r/R\approx 2$ in all directions except vertically, where particle pair depletion is observed. Beyond this region, the function displays complex interwoven peaks and minima, reflecting a complex microstructure arrangement. In contrast, the pair distribution function for cubes shows a stronger ordering in the horizontal direction at high density. Distinct vertical lines appear at distances around $r/R=1.2$ to $1.4$, followed by separations at $r/R=1.7$ to $1.9$ and $r/R=2.4$ to $2.9$, indicating a preferred arrangement of cubes in the $xy$-plane in a square lattice with a lattice spacing $d\approx 1.3$. The continuous lines along the $z$-axis indicate that cubic particles lack ordering along the gravity direction. This suggests that cubes form columnar structures capable of sliding past one another within the densely packed suspension (Fig. 8).
@@ 3.6 Segregation of spherical/cubic particles in a bidisperse suspension
Gravitational sedimentation is a fundamental technique for solid–liquid separation, widely applied in polydisperse systems where size segregation and particle sorting are essential. Although monodisperse systems, comprised of particles of uniform size, have been extensively explored [54, 55, 21, 56, 57, 12], the dynamics of polydisperse suspensions (systems with particles of varying sizes, densities, or shapes) remain a topic of significant interest. Polydisperse systems present complex behavior, as the distribution of particle sizes and densities critically influences the overall flow response. Predicting the settling velocity of each particle class in a polydisperse suspension is a primary challenge; understanding these velocities is essential, as they shape the concentration profile and enable effective particle separation by size. Here, we focus on bidisperse sedimentation, which serves as a simplified yet valuable model for examining inter-particle interactions within polydisperse systems. In bidisperse suspensions with low total volume fractions, the hindrance effect becomes more complex than in monodisperse suspensions due to differential impacts of fluid displacement on the settling velocities of distinct particle classes. Although previous studies have measured settling velocities in bidisperse suspensions, these primarily focus on the settling velocities of interfaces or isolated particles within distinct zones of the suspension. When translational symmetry in the vertical direction is disrupted, for example, by the presence of horizontal boundaries, distinct concentration regions can form, leading to varied settling dynamics. In such cases, measurements are typically concentrated on the settling velocities at the interfaces between these zones [58, 59, 60, 61]. In bidisperse suspensions, where particles differ in radii, two primary sedimentation fronts generally propagate from top to bottom: one at the boundary between clear fluid and small particles, and another between small particles and a homogeneous suspension containing both particle sizes. The ratio of particle densities and sizes significantly affects size-based segregation in these systems. Numerical simulations using Stokesian dynamics have been conducted on bidisperse suspensions, covering size ratios up to 4 [20, 62]. Recently, Li and Botto [63] examined the settling velocity statistics for dilute, non-Brownian, polydisperse sphere suspensions using Stokesian Dynamics, analyzing how total volume fraction and particle size distribution width influence sedimentation behavior.
@@
In this section, we examine the settling velocity of non-Brownian bidisperse suspensions composed of a mixture of cubes and spheres, spanning a range of particle volume fractions from low to high. We assume all particles have a uniform mass density and identical volumes. To achieve this, we calculate an equivalent side length for the cubes, ensuring that each cube has the same volume as a sphere with a unit radius. In the bidisperse case, we consider a total of $150$ particles, evenly divided between the two particle shapes. We define the total solid volume fraction ($\phi$) as the sum of the volume fractions of each particle type: $\phi=\phi_{1}+\phi_{2}$, where $\phi_{1}$ and $\phi_{2}$ represent the volume fractions of spheres and cubes, respectively. We examine the case of equal volume fractions for both particle types, i.e., $\phi_{1}=\phi_{2}$. Additionally, an external force of magnitude $F_{b}=9.07\,$\mathrm{p}\mathrm{N}$$ is exerted along the z-axis to represent the gravitational pull. As illustrated in Fig. 9, the sedimentation velocity of the bidisperse suspension is strongly influenced by the volume fraction of each species. To reduce statistical error, the velocity distributions are averaged across several random initial configurations of particles within a triply periodic simulation box. In the most dilute case ($\phi=0.03$), the suspension requires a longer time to reach a steady state. The average sedimentation velocity of each species plays a key role in the overall settling dynamics of the suspension. As shown in Fig. 9a, spheres settle approximately $13.2\%$ faster than cubes. This disparity is due in part to the higher angular rotation induced by the sharp edges of the cubes [36, 37], which enhances the hindrance effect relative to that experienced by spheres. As the volume fraction increases, interparticle hydrodynamic interactions become more pronounced, causing a moderate reduction in the settling velocity of each species. At higher volume fractions ($\phi=0.36$), the suspension quickly reaches a steady state, with the reduced settling velocity attributed to increased hindrance effects and diminished velocity fluctuations in both perpendicular and parallel directions. The study of Li and Botto [63] demonstrated that in polydisperse mixtures of small and large spheres, the normalized average settling velocity decreases more rapidly with increasing $\phi$ for smaller particles. Our study confirms that spheres settle faster than cubes at all volume fractions, with segregation efficiency improving at lower $\phi$.
@@
Figure 10: (a) The average settling velocity of the particles, $U_{\text{sed}}$, normalized by the Stokes velocity, $U_{0}$, is shown as a function of volume fraction ($\phi$). In this context, $U_{0}$ represents the settling velocity of an isolated sphere in the same periodic domain. Dimensionless hydrodynamic velocity fluctuations of particles as a function of volume fraction ($\phi$) are shown for (b) the settling direction ($\sigma_{\parallel}$) and (c) the direction perpendicular ($\sigma_{\perp}$) to it.
Fig. 10 presents the normalized average settling velocity and the corresponding velocity fluctuations for different particle types in a bidisperse mixture across various bulk concentrations. The hindered settling behavior follows a trend similar to that observed in monodisperse suspensions (Fig. 4), however, the velocity fluctuations show significant difference for the two species. For dilute bidisperse suspensions (i.e., $\phi<0.1$), Fig. 10a shows a notable distinction in the sedimentation velocity between two species. In this regime, velocity fluctuations in both the vertical ($\sigma_{\parallel}$) and horizontal ($\sigma_{\perp}$) directions are pronounced for both particle types compared to those in monodisperse suspensions, resulting in a lower overall settling velocity for the bidisperse mixture (Fig. 10b,c). As the solid volume fraction increases, the interparticle distances diminish, leading to more frequent collisions and stronger hydrodynamic interactions. This effect causes a peak in vertical velocity fluctuations ($\sigma_{\parallel}$) around $\phi\approx 0.23$ for cubic particles. In the range $0.03\leq\phi\leq 0.23$, vertical velocity fluctuations for cubes rise from $12.5\%$ to $40\%$ compared to those of spheres, while fluctuations in the horizontal direction exhibit a steady increase over this range. The forces induced by higher rotation in cubes result in more pronounced horizontal dispersion compared to spheres. Further increment in $\phi$, suppresses velocity fluctuations in both directions due to augmented hydrodynamic interactions, which stabilize the motion and reduce overall fluctuations.
@@
Figure 11: The probability density function (PDF) of vertical velocity is presented for suspensions of (a) spherical and (b) cubic particles across different solid volume fractions. Negative values of $U_{i}/U_{\text{sed}}$ indicate that particles are moving vertically upward, against the direction of gravity. Additionally, (c) illustrates the percentage of particles moving opposite to the gravity at any given moment as a function of solid volume fractions.
In the previous section, we discussed the overall settling behavior and velocity fluctuations of the particle mixture. Here, we examine the normalized probability density function (PDF) of particle velocities in the vertical direction (Fig. 11). These PDFs were generated by considering all particles within the simulation domain. Although the PDF for bidisperse suspensions follows a similar pattern to that of monodisperse suspensions (Fig. 6), there is a notable distinction in the distribution of vertical velocities between the two species in a bidisperse mixture. Specifically, the longer tail in the PDF of vertical velocity for cubic particles indicates that a significant fraction of cubes exhibit upward motion (against gravity) compared to spheres, particularly at moderate volume fractions ($0.08\leq\phi\leq 0.25$) (Fig. 11a,b). As described in Sec. 3.4, we also calculate the percentage of upward-moving particles ($N_{U_{z}}$), as shown in Fig. 11c. In the range $0.02\leq\phi\leq 0.13$, $N_{U_{z}}$ for cubes is nearly double that of spheres, indicating that disturbances created by surrounding spheres significantly influence the upward movement of cubes. This behavior arises from higher velocity fluctuations in both vertical and horizontal directions for cubes within this moderate range, leading to greater particle dispersion in the plane perpendicular to gravity. At higher volume fractions, increased collisions and intensified hydrodynamic interactions further hinder settling velocities, causing a larger proportion of particles (both spheres and cubes) to fall with velocities close to the mean suspension velocity. Consequently, only $3.5\%$ of cubic particles and $2.6\%$ of spherical particles exhibit upward motion at higher $\phi$, reflecting the impact of enhanced particle interactions on settling dynamics.

### v2

@@ 3.5 Probability density function of particle sedimentation velocity
To better understand the differences in velocity fluctuations between spherical and cubic particles, we investigate the instantaneous particle velocity during sedimentation. Figs 7a and 7b illustrate the normalized probability density function (PDF) of particle velocity as a function of bulk concentration. Notably, $U_{i}/U_{\text{sed}}>1$ indicates that particles are falling faster than the average sedimentation velocity. Positive values of $U_{i}/U_{\text{sed}}$ imply downward movement (i.e. in the direction of gravity), while negative values represent upward movement. Note that as the particles settle, the surrounding fluid generates an upward jet due to the conservation of mass, that can drag some particles with it. Interesting, the variance of the velocity in the vertical direction is a non-monotonous function of the volume fraction. In a very dilute regime ($\phi\approx 0.001$), both the spherical and cubic suspensions exhibit a sharp peak, indicating that the velocity variance is small. In this range, the larger inter-particle separation distance results in weaker hydrodynamic interactions, thus deviation from the single particle dynamics are small. Additionally, the percentage of upward-moving particles ($N_{U_{z}}$) is calculated by integrating the area under the PDF of the settling velocity. Figure 7c illustrates that, in the highly dilute regime, the fraction of upward-moving particles approaches to nearly zero for both types of particles.
@@
As the volume fraction increases, velocity fluctuations increase, causing particles to disperse more and collide more frequently. At moderate concentrations, hydrodynamic interactions among nearby settling particles and inter-particle collisions create a strong upward jet, which can lift some particles (both spheres and cubes) along with it. The frequent hydrodynamic interactions and particle rebounds after collisions may also lead to individual particles moving upward. A closer examination of the PDF of vertical particle velocity reveals that the tails become stretched as $\phi$ increases from $0.025$ to $0.125$, with the left tail slightly longer than the right. Therefore, the fraction of upward-moving particles increases with rising bulk concentration in the low to moderate range. Results in Fig. 7c indicate that in cubic suspensions, approximately $5.3\pm 0.2\%$ of particles move against the settling direction, compared to only $4.6\pm 0.1\%$ for spherical particles at $\phi\approx 0.1$. In this range, a closer inspection in Figs. 7a,b reveals that the peak of the PDF is slightly higher for spherical particles, indicating that a greater number of them attain velocities close to the mean sedimentation velocity. In contrast, cubic particles exhibit a slightly stretched left tail in the negative region, showing a broader distribution of lower velocities, see also Supplementary Figure 1. The sharp edges and rapid changes in hydrodynamic forces due to variations in orientation make cubes more susceptible to strong translational and rotational displacements, as well as velocity fluctuations. Additionally, cubes tend to acquire slightly higher angular velocities than spheres, resulting in significant rotation-induced forces perpendicular to their direction of movement. Consequently, cubic suspensions generate larger rotation-induced forces due to the higher magnitude of angular rotation, resulting in a greater fraction of particles being entrained by the upward jet compared to spherical suspensions. At moderate $\phi$, this effect becomes more pronounced, with $6.2\pm 0.2\%$ of cubic particles moving upward, compared to only $2.86\pm 0.08\%$ for spherical particles at $\phi\approx 0.24$ (Fig. 7c).
However, as $\phi$ increases further, the system becomes more densely packed, suppressing velocity fluctuations in both directions. Consequently, the tails of the PDF begin to contract for both particle types (Fig. 7a,b), leading to a gradual decrease in the percentage of upward-moving particles, as shown in Fig. 7c. In the very dense regime ($\phi=0.5$), only a small fraction of spherical particles ($0.070\pm 0.003\%$) exhibit instantaneous upward movement, whereas a slightly larger fraction of cubic particles ($1.3\pm 0.1\%$) continue to move upward.
@@ 3.6 Suspension microstructure
For dilute suspensions ($\phi\simeq 2.5\%$), the radial distribution function (RDF) exhibits a subtle peak for both spherical and cubic particles, though it is slightly lower for cubic suspensions (Fig. 8). In this regime, weak hydrodynamic interactions allow particles to form transient pairs, maintaining contact for extended periods. Consequently, a slightly higher local density of spherical particles is observed around the reference particle compared to cubic particles in this regime. As $\phi$ increases, hydrodynamic interactions among multiple particles become more pronounced, leading to the formation of aggregates or clusters in close contact, as reflected in the sharper RDF peaks. Unlike spherical particles, which maintain a fixed contact point at $r/R=2$ (Fig. 8a), cubic particles interact over a broader range of center-to-center distances [36]. As a result, the RDF profile for cubic particles appears to be more oblique (Fig. 8b). At a moderate solid volume fraction ($\phi\simeq 36\%$), the particle-pair distribution undergoes a gradual transition. Initially, particle pairs accumulate around $r/R=2.2$ for spheres and $r/R=2.33$ for cubes. This enrichment later shifts to a depletion at greater distances, approximately $r/R\simeq 3.4$ for spheres and $r/R\simeq 3.5$ for cubes. The reduced inter-particle spacing compared to dilute conditions increases the likelihood of cluster formation at short distances. This effect is more pronounced for spherical particles, as evidenced by their higher RDF peaks. In contrast, cubic particles exhibit a greater tendency to escape clustering due to their higher rotational velocities, a behavior also reported by Seyed-Ahmadi and Wachs [36]. With further increment in $\phi$, intensified hindered settling results from reduced inter-particle distances, leading to a more isotropic particle structure. This is reflected in the persistent oscillations beyond the first RDF peak for both suspensions at $\phi\simeq 0.5$. At larger distances from the reference particle ($r/R\geq 5.1$ for both cases), the particle distribution becomes random, with $g(r)\simeq 1$.
@@
As the volume fraction increases, short-range hydrodynamic interactions between nearby particles become more pronounced, causing more particle pairs to cluster around the reference particle and enhancing structural ordering. We observe a transition from a preferential horizontal alignment at lower bulk concentrations ($0.025\leq\phi\leq 0.125$) to a more concentric, ring-like distribution at moderate concentrations ($0.15\leq\phi\leq 0.4$). A spherical contour of particle pairs forms at distances of $r/R=2.05$ to $2.65$ for spheres and $r/R=1.15$ to $1.6$ for cubes at $\phi=0.36$ (Fig. 9) reveling an isotropic ordering. At this volume fraction a second ring due to second neighbors becomes visible.
The most intriguing case occurs at the highest volume fraction studied, $\phi=0.5$, where the pair distribution functions of cubes and spheres exhibit striking differences. For spheres, the pair distribution function shows a high concentration of particles at $r/R\approx 2$ in all directions. Beyond this region, the pair distribution function shows strong fluctuations. As the system is very crowded spheres take longer to explore the configurational space and it is difficult to reduce the statistical errors. In contrast, the pair distribution function for cubes shows a stronger ordering in the horizontal direction at high density. Distinct vertical lines appear at distances around $r/R\approx 2$, followed by separations at $r/R\approx 2.7$ and $r/R\approx 4.5$, indicating a preferred arrangement of cubes in the $xy$-plane in a square lattice with a lattice spacing $d\approx 2R\approx 1.25h$, where $h=\left(4\pi/3\right)^{1/3}R$ is the cube side length. The continuous lines along the $z$-axis indicate that cubic particles lack ordering along the gravity direction. This suggests that cubes form columnar structures capable of sliding past one another within the densely packed suspension, see also Supplementary Movie 1.
@@ 3.7 Segregation of spherical/cubic particles in a bidisperse suspension
Gravitational sedimentation is a fundamental technique for solid–liquid separation, widely applied in polydisperse systems where size segregation and particle sorting are essential. Although monodisperse systems, comprised of particles of uniform size, have been extensively explored [56, 57, 21, 58, 59, 12], the dynamics of polydisperse suspensions (systems with particles of varying sizes, densities, or shapes) remain a topic of significant interest. Polydisperse systems present complex behavior, as the distribution of particle sizes and densities critically influences the overall flow response. Predicting the settling velocity of each particle class in a polydisperse suspension is a primary challenge; understanding these velocities is essential, as they shape the concentration profile and enable effective particle separation by size. Here, we focus on bidisperse sedimentation, which serves as a simplified yet valuable model for examining inter-particle interactions within polydisperse systems. In bidisperse suspensions with low total volume fractions, the hindrance effect becomes more complex than in monodisperse suspensions due to differential impacts of fluid displacement on the settling velocities of distinct particle classes. Although previous studies have measured settling velocities in bidisperse suspensions, these primarily focus on the settling velocities of interfaces or isolated particles within distinct zones of the suspension. When translational symmetry in the vertical direction is disrupted, for example, by the presence of horizontal boundaries, distinct concentration regions can form, leading to varied settling dynamics. In such cases, measurements are typically concentrated on the settling velocities at the interfaces between these zones [60, 61, 62, 63]. In bidisperse suspensions, where particles differ in radii, two primary sedimentation fronts generally propagate from top to bottom: one at the boundary between clear fluid and small particles, and another between small particles and a homogeneous suspension containing both particle sizes. The ratio of particle densities and sizes significantly affects size-based segregation in these systems. Numerical simulations using Stokesian dynamics have been conducted on bidisperse suspensions, covering size ratios up to 4 [20, 64]. Recently, Li and Botto [65] examined the settling velocity statistics for dilute, non-Brownian, polydisperse sphere suspensions using Stokesian Dynamics, analyzing how total volume fraction and particle size distribution width influence sedimentation behavior. They also reported that in a polydisperse mixtures of small and large spheres, the normalized average settling velocity decreases more rapidly with increasing $\phi$ for smaller particles.
@@
In this section, we examine the settling velocity of non-Brownian bidisperse suspensions composed of a mixture of cubes and spheres, spanning a range of particle volume fractions from low to high. We assume all particles have a uniform mass density and identical volumes. To achieve this, we calculate an equivalent side length for the cubes, ensuring that each cube has the same volume as a sphere with a unit radius. In the bidisperse case, we consider a total of $150$ particles, evenly divided between the two particle shapes. We define the total solid volume fraction ($\phi$) as the sum of the volume fractions of each particle type: $\phi=\phi_{1}+\phi_{2}$, where $\phi_{1}$ and $\phi_{2}$ represent the volume fractions of spheres and cubes, respectively. We examine the case of equal volume fractions for both particle types, i.e., $\phi_{1}=\phi_{2}$, and apply equal downward force ($f=1$) on all the particles. As illustrated in Fig. 10, the sedimentation velocity of the bidisperse suspension is strongly influenced by the volume fraction of each species. To reduce statistical error, the velocity distributions are averaged across several random initial configurations of particles within a triply periodic simulation box. In the most dilute case ($\phi=0.03$), the suspension requires a longer time to reach a steady state. The average sedimentation velocity of each species plays a key role in the overall settling dynamics of the suspension. As shown in Fig. 10a, spheres settle faster than cubes. This disparity is due in part to the higher angular rotation induced by the sharp edges of the cubes [36, 37], which enhances the hindrance effect relative to that experienced by spheres. As the volume fraction increases, interparticle hydrodynamic interactions become more pronounced, causing a moderate reduction in the settling velocity of each species. At higher volume fractions ($\phi=0.36$), the suspension quickly reaches a steady state, with the reduced settling velocity attributed to increased hindrance effects and diminished velocity fluctuations in both perpendicular and parallel directions. Our study shows that spheres settle faster than cubes at all volume fractions, with segregation efficiency improving at lower $\phi$.
@@
Figure 11: The average settling velocity of the particles, $U_{\text{sed}}$, normalized by the Stokes velocity, $U_{0}$, is shown as a function of volume fraction ($\phi$). In this context, $U_{0}$ represents the settling velocity of an isolated sphere in the same periodic domain. Here, the error bars are smaller than the size of the symbols. Inset shows the difference of mean sedimentation velocity of the bisdisperse suspension with that of a monodisperse suspension of spheres.
Fig. 11 presents the normalized average settling velocity for different particle types in a bidisperse mixture across various bulk concentrations. The hindered settling behavior follows a trend similar to that observed in monodisperse suspensions (Fig. 5), but the mean settling velocity of individual species differs significantly. While the overall velocity of the mixture is slightly lower than that of a monodisperse suspension of spheres in the dilute limit, the variation in settling velocities between species is notable. Fig. 11 shows that spheres settle at a rate of $11\%$ faster than cubes in this range (i.e. $U^{\text{sphere}}_{\text{sed}}/U^{\text{cube}}_{\text{sed}}=1.11$). In a mixture of regular and irregular particles, spheres exhibit the most hydrodynamic shape, minimizing drag and enabling faster settling. In contrast, irregular particles like cubes, with sharp edges and flat faces, experience higher drag, leading to slower descent. As the volume fraction ($\phi$) increases the generated upward flux of the surrounding fluid increases which slows the settling rate. In this regime the relative velocity difference between spheres and cubes increases to $34\%$ at $\phi\approx 0.25$. With further increases in $\phi$, hindered settling becomes more pronounced, reducing somewhat the segregation efficiency. At higher concentrations ($\phi\approx 0.36$), the relative velocity difference narrows to $24\%$, and the overall mixture velocity approaches to a comparable magnitude that of a monodisperse suspension of spheres.
Figure 12: Dimensionless hydrodynamic velocity fluctuations of particles as a function of volume fraction ($\phi$) are shown for (a) the settling direction ($\sigma_{\parallel}$) and (b) the direction perpendicular ($\sigma_{\perp}$) to it. (c) Comparison of anisotropy in velocity fluctuations of bidisperse mixture with the monodisperse sphere suspension.
In a polydisperse suspension the particles shape also affects the velocity fluctuations. Fig. 12a,b presents the velocity fluctuations of individual species and the entire mixture in directions parallel and perpendicular to gravity. For dilute bidisperse suspensions ($\phi\approx 0.1$), Fig. 12a shows a clear difference in the velocity fluctuations in the vertical direction between the two species. In this regime, velocity fluctuations in both directions are more pronounced than in monodisperse suspensions. Specifically, the fluctuations of the cube velocities in the vertical direction are a $26\%$ higher in the mixture than in the monodisperse suspension while for spheres the relative difference it is only a $4\%$. This results in an overall rise in vertical velocity fluctuations in the bidisperse mixture compared to monodisperse cases.
As the solid volume fraction increases, reduced interparticle distances lead to more frequent collisions and stronger hydrodynamic interactions. This effect causes a rise in vertical velocity fluctuations ($\sigma_{\parallel}$) around $\phi\approx 0.23$ for both particle types. In this range, cubic particles in a bidisperse mixture show a $40\%$ increase in $\sigma_{\parallel}$ compared to monodisperse cubic suspensions, while spherical particles exhibit a $27\%$ increase. Simultaneously, horizontal velocity fluctuations steadily increase for both species. These trends collectively alter the degree of anisotropy in velocity fluctuations, as shown in Fig 12c. The combined contributions of individual species lead to a net rise in both $\sigma_{\parallel}$ and $\sigma_{\perp}$ for the entire mixture, directly influencing the anisotropy of velocity fluctuations. As the volume fraction increases from low to moderate values, the anisotropy parameter, $\Gamma$ decreases (Fig. 12c). While $\Gamma$ follows the same qualitative trend as in Fig 6c, a noticeable difference emerges between the whole mixture and monodisperse sphere suspensions. However, the absolute difference diminishes as the solid volume fraction increases due to a steady rise in horizontal fluctuations. At higher $\phi$, intensified hydrodynamic interactions suppress velocity fluctuations in both directions, stabilizing particle motion and leading to a gradual decay in $\sigma_{\parallel}$ and $\sigma_{\perp}$ for both species. Consequently, the overall anisotropy of the bidisperse mixture decreases, though it remains slightly higher than in the monodisperse case. The pair distribution function shows a weaker ordering in the bidisperse suspension compared with the monodisperse suspensions, see Supplementary Fig. 2. The weaker ordering could explain the higher velocity fluctuations observed in bidisperse suspensions.
@@
Figure 13: The probability density function (PDF) of the vertical velocity for (a) spherical and (b) cubic particles in a bidisperse suspension is shown across different solid volume fractions. Negative values of $U_{i}/U_{\text{sed}}$ indicate particles moving upward, opposing gravity. Additionally, (c) depicts the percentage of particles ($N_{U_{z}}$) moving against gravity at any given moment as a function of solid volume fraction. Inset of (c) compares $N_{U_{z}}$ in a bidisperse suspension to that in a monodisperse system.
In addition to examining the overall settling behavior and velocity fluctuations of the particle mixture, we also analyze the normalized probability density function (PDF) of vertical particle velocities (Fig. 13) for all particles within the simulation domain. While the PDF for bidisperse suspensions follows a trend similar to that of monodisperse suspensions (Fig. 7), a key distinction emerges in the distribution of vertical velocities between the two particle species in a bidisperse mixture. Across all solid volume fractions, the PDF of vertical velocity exhibits a slightly higher peak for spherical particles than for cubic particles, indicating that a larger fraction of spheres settle with velocities close to their mean sedimentation velocity. Additionally, a small fraction of particles moves instantaneously against gravity (see Supplementary Movie 2), even in the dilute regime ($\phi<0.1$; Fig. 13a,b). The longer tail in the PDF for cubic particles suggests that cubes are more likely to exhibit upward motion than spheres, contributing to the overall velocity distribution of the mixture. As shown in Fig. 13c, $7.67\pm 0.27\%$ of cubic particles display upward motion, compared to $4.94\pm 0.2\%$ for spherical particles. This results in a total of $6.38\pm 0.23\%$ of particles in the bidisperse suspension moving instantaneously against gravity, higher than the $5.27\pm 0.17\%$ observed in monodisperse suspensions (Inset of Fig. 13c).
At moderate volume fractions, the tail of the PDF extends further into the negative velocity region due to increased velocity fluctuations and particle dispersion in the plane perpendicular to gravity. Consequently, more particles exhibit upward motion in bidisperse suspensions than in monodisperse ones. As shown in the inset of Fig. 13c, at $\phi=0.23$, $7.39\pm 0.2\%$ of particles in the bidisperse suspension move upward, compared to $6.14\pm 0.17\%$ in the monodisperse case (see also Supplementary Movies 3 and 4). However, as $\phi$ increases further, intensified hydrodynamic interactions and inter-particle collisions suppress velocity fluctuations in both directions. As a result, at $\phi=0.36$, only $6\pm 0.26\%$ of cubic particles and $3\pm 0.14\%$ of spherical particles exhibit upward motion, highlighting the impact of enhanced particle interactions on the settling dynamics of the mixture.